# Information Discovery in e-Commerce

### v1

@@ 7.2.3 Knowledge-grounded dialogue systems
Although answering inquiries is essential for dialogue systems, especially for task-oriented dialogue systems, it is still far behind a natural knowledge-grounded dialogue system, which should be able to understand the facts involved in the current dialogue session (so-called facts matching) and diffuse them to other similar entities for knowledge-based dialogues (i.e., entity diffusion):
@@
    Facts matching: In dialogue systems, matching utterances to exact facts is much harder than explicit factoid inquiries answering. Though some utterances, whose subjects and relations can be easily recognized, are fact-related inquiries for some utterances, the subjects and relations are elusive, leading to trouble matching exact facts. Table 7.1 shows an example of Items 1 and 2 talking about the film “Titanic”. Unlike item 1, which is a typical question-answering conversation, item 2 is a knowledge-related chit-chat without any explicit relation. It is difficult to define the exact fact match for item 2.
@@
    Entity diffusion: Another noticeable phenomenon is that the conversation usually drifts from one entity to another. In Table 7.1, utterances in items 3 and 4 are about the entity “Titanic.” However, the entity of responses is other similar films. The current knowledge triplets rarely capture such entity diffusion relations. The response in item 3 shows that the two entities, “Titanic” and “Waterloo Bridge”, are relevant through “love stories”. Item 4 suggests another similar shipwreck film “Titanic”.
Knowledge-grounded dialogue (KGD) systems are proposed to address the aforementioned challenges. According to the knowledge modality, existing works for knowledge-grounded dialogue systems can be mainly categorized into two groups. The methods in the first group aims to leverage structured knowledge (given knowledge graphs) [wu2020diverse, zhou-etal-2020-kdconv, wang2020improving, wu2020improving, xu2020knowledge, xu2020conversational, jung2020attnio, tuan2019dykgchat, wu2019proactive, moon2019opendialkg, zhou2018commonsense, liu2018knowledge]; while the ones in the second group focus on leveraging unstructured knowledge, such as document-based unstructured knowledge (given a whole document, e.g. Wikipedia article) [chuanmeng2020refnet, ma2020survey, ma2020compare, tian-etal-2020-response, ren2019thinking, gopalakrishnan2019topical, zekangli2018incremental, qin-etal-2019-conversing, moghe2018towards, zhou2018dataset, parthasarathi2018extending] or piece-based unstructured knowledge (given some separate pieces of knowledge, e.g. Foursquare tips) [ghazvininejad2018knowledge, dinan2018wizard, meng2020dukenet, kim2020sequential, Lian2019Learning, zheng2020difference, chen2020bridging, zheng2020approximation, zhao2020knowledge, zheng2019enhancing, lin-etal-2020-generating]. There are key research directions for both groups: 1. improving knowledge selection [kim2020sequential]; 2. improving knowledge-aware response generation [zhao2020knowledge] or response selection [young2018augmenting, zhao2019document, hua2020learning, sun2020history]; that is, given the chosen knowledge, how to better generate a response token by token or select a response from pre-defined response candidates; 3. leveraging multiple knowledge modalities [liu2019knowledge, zhao2020multiple]; that is, how to utilize structured, unstructured, and even other types of knowledge simultaneously; 4. overcoming data scarcity [zhao2019low, li2020zero].
@@ 4 A: Is there anything like the Titanic?
The neural knowledge diffusion (NKD) model introduces knowledge into the dialogue generation. This method can match the relevant facts for the input utterance and diffuse them to similar entities [liu2018knowledge].
@@
Early studies on knowledge selection in KGD calculated the weight of each piece of knowledge and got the weighted sum of their representations [ghazvininejad2018knowledge, zheng2019enhancing, lin-etal-2020-generating, zheng2020approximation]. bordes2016learning employed memory networks to handle restaurant reservations, using a small number of keywords to handle entity types in a knowledge base (cuisine type, location, price range, party size, rating, phone number, and address). ghazvininejad2017knowledge adapt it to memorize the relevant, grounded facts for a neural conversation model. The hierarchical variational memory network (HVMN) adds the hierarchical structure and the variational memory network into a neural encoder-decoder network for non-task-oriented dialogue generation [ziwwsdm1].
@@
Several recent studies on knowledge selection focused on calculating a weight on each piece of knowledge and then directly sampling the amount of knowledge with the highest weight. Specifically, dinan2018wizard proposed the TMemNet model that uses context to predict a distribution over pieces of knowledge and then only sample one of them into a decoder. They also introduced a knowledge selection loss to supervise knowledge selection during training. Lian2019Learning proposed PostKS, which uses a context to predict a prior distribution over pieces of knowledge. During training, the prior distribution is supervised by a posterior distribution predicted by the context and the corresponding response. Similar to PostKS, zheng2020approximation proposed to use a context and a piece of knowledge retrieved by the context to predict a distribution over fragments of knowledge, where the probability of the amount of knowledge retrieved by the corresponding response is maximized during training. The former distills a context containing multiple utterances at different turns into a vector that is used to match with the representation of a piece of knowledge to get a score, while the latter matches every utterance in a context with a piece of knowledge to get matching features that are aggregated to get a score. A piece of knowledge is chosen based on the score list for all pieces of knowledge. kim2020sequential proposed a sequential knowledge transformer (SKT), which jointly uses previously selected knowledge and context to facilitate knowledge selection. chen2020bridging upgraded SKT by adding the knowledge distillation-based training strategy to improve knowledge selection. zheng2020difference proposed a method that introduces the difference information between the previously selected knowledge and the current pieces of candidate knowledge to facilitate knowledge selection. meng2020dukenet proposed DukeNet, which regards tracking the previously selected knowledge and selecting the current knowledge as dual tasks within a dual learning paradigm [qin2020dual]. zhao2020knowledge proposed an RLKS method, where the selected knowledge is sent to a decoder to generate a response that would be compared with the ground truth response to give feedback to further supervised knowledge selection. meng2021initiative proposed a mixed-initiative knowledge selection method for KGC, which explicitly distinguishes between user-initiative and system-initiative knowledge selection at each conversation turn to improve the performance of knowledge selection. sun2021conversations find that the amount of knowledge available in different languages is highly imbalanced. Hence the authors proposed the cross-lingual knowledge grounded conversation with a self-distillation knowledge selection and curriculum learning.
@@
Recent years have witnessed a rapid development of pre-trained language models in open-domain dialogue systems. Large pre-trained language models can store knowledge into their parameters during pre-training and can generate informative responses in conversations [Zhao2020ArePL]. Petroni2019LanguageMA have shown that pre-trained language models can serve as knowledge bases for downstream tasks (e.g., question-answering [Roberts2020HowMK]). On this basis, Zhao2020ArePL have shown that pre-trained language models can ground open-domain dialogues using their implicit knowledge. madotto2020learning embeded knowledge bases into model’s parameters for end-to-end task-oriented dialogues. Roller2021RecipesFB fine-tuned pre-trained language models on KGD data. Cui2021KnowledgeEF proposed knowledge-enhanced fine-tuning methods to handle unseen entities. Xu2021RetrievalFreeKD proposed a topic-aware adapter to adapt pre-trained language models in KGDs. Liu2022MultiStagePF proposed a multi-stage prompting approach for triggering knowledge in pre-trained language models. Wu2022LexicalKI proposed lexical knowledge internalization to integrate token-level knowledge into the model’s parameters. The problem of hallucination is becoming more and more challenging. sun2023contrastive optimized the implicit knowledge eliciting process, i.e., reduce hallucination of pre-trained language models in KGD, via a contrastive learning framework.
@@ 7.2.4 Empathetic dialogue generation
With the rise of data-driven learning approaches [SutskeverVL14, vaswani2017attention], open-domain dialogue generation models have seen increasing interest in recent years [vinyals2015neural, shang, SerbanLCP16, limrjgg16, zhou2018commonsense, dinan2018wizard].
Recent approaches generate emotional responses conditioning based on a manually specified label to control the dynamic content of the target output [zhou2018emotional, lis18, zhou2018mojitalk, huang2018automatic, wei2019emotion, colombo2019affect, shenf20].
To control the emotional content of the target output, recent approaches generate emotional responses conditioning on a manually specified label [zhou2018emotional, lis18, zhou2018mojitalk, huang2018automatic, wei2019emotion, colombo2019affect, shenf20].
Unlike emotional dialogue generation, the study of empathetic dialogue generation avoids an additional step of determining which emotion type to respond to explicitly [skowron2013affect]. Several studies [Rashkin18, zhong2019affect, Shin19, rashkin2019towards, santhanam19, linmsxf19, lin2020caire, zhong2020, MajumderHPLGGMP20, li2020empdg] have attempted to make dialogue models more empathetic. rashkin2019towards combined existing models in different ways to produce empathetic responses. While linmsxf19 softly combined the possible emotional responses from several separate experts. MajumderHPLGGMP20 considered of this polarity-based emotion clusters and emotional mimicry. li2020empdg proposed a multi-resolution adversarial framework that considers multi-granularity emotion factors and users’ feedback. li2020empathetic investigated how to leverage external knowledge to explicitly improve the emotional understanding and expression in the task of empathetic dialogue generation. sabour2022cem focused on two aspects in empathetic dialogue generation: affection and cognition. The authors proposed a method with various commonsense reasoning to improve understanding of interlocutors’ situations and feelings.
Besides the advancements in empathetic dialogue models, the emergence of new emotion-labeled dialogue corpora has also contributed to this research field [li2017dailydialog, hsuckhk18, rashkin2019towards]. rashkin2019towards considered a more prosperous and evenly distributed set of emotions and released a dataset EmpatheticDialogues, where a listener responds to a speaker in an emotional situation in an empathetic way.
@@ 7.3 Emerging directions
This section describes recent emerging directions on question-answering and dialogue systems in e-commerce. These emerging directions on QA and dialogue systems can be divided into $5$ perspectives: safety, ethics, interpretability, privacy, and evaluation.
@@
As discussed in Section 7.1, e-commerce QA agents aim to answer these questions through massive reviews. However, reviews may not answer these questions as they may not contain any relevant answers for the question, or a query may be poorly phrased and therefore requires additional clarification. Moreover, untruthful comments and spams are also widely observed in e-commerce reviews. These answers or reviews may attribute to multiple factors such as misunderstandings of the question, improper expressions during writing, and even malicious attacks from the competitors [carmel2018product]. mihaylova2019semeval investigated the fact-checking problem in a QA scenario with a system to classify the veracity of answers. zhang2020answerfact released a large-scale fact-checking dataset called AnswerFact for investigating the answer veracity in e-commerce QA. estes2022fact developed a high-speed fact-verification system that has a very high false statement recall and very high true statement precision to product question-answering. However, the authors still applied a rule-based method in their system, and found that pre-trained language models are still unable to perform fact-checking well on the e-commerce structured catalog data. As limitations such as poor generalization exist in rule-based methods, how to optimize pre-trained language models in fact-checking for product question-answering still needs more attention in future research.
@@
In dialogue systems, ethical challenges have been becoming more and more severe in recent years. Currently, most dialogue systems are developed from scratch with large corpora or fine-tuned through pre-trained language models. Large-scale datasets collected from the open Internet have been applied during model training. However, offensive and malevolent content can be observed in the data [si2022so]. To avoid being unintentionally offensive or harming the user, research studies have been performed to detect toxic speech around religion, race, and violence [tripathi2019detecting, dinan2020queens, zhang2021human, kann2022open, si2022so]. zhang2021human proposed a human-machine collaborative evaluation framework for reliable toxic speech detection in dialogue systems. Note that si2022so studied toxic speech in open-domain dialogue systems to reveal that specific kinds of “non-toxic” queries are able to trigger an open-domain conversation artificial intelligence to output toxic responses. However, how to respond when these malevolence topics are being identified is still an open question for conversational artificial intelligence [kann2022open]. As more and more e-commerce dialogue systems have also been trained based on pre-trained language models, ethical challenges will still need to be tackled in future research.
@@
Poor explainability is a challenging problem for most e-commerce QA approaches. Most existing e-commerce QA approaches apply end-to-end semantic matching methodologies, which tend to be black-box and directly output a matching score for each question answer pair. zhao2019riker address the explainable QA problem through a hybrid retrieval-based framework. The authors employ a bidirectional recurrent neural network in the internal word representation stage and apply a keyword-aware retrieval method during the second stage. In contrast, the tf-idf ranking function naturally exhibits much better interpretability owing to its transparency and intuitiveness. The conversational recommendation is another typical application of e-commerce dialogue systems [sun2016conversational, mangili2020bayesian]. Most approaches neglect explainability during the recommendation action learning procedure. As the first attempt, chen2020towards propose an incremental multi-task learning framework using user feedback for the task of explainable conversational recommendation. By considering user preferences as latent variables in a variational Bayesian manner, ren2022variational employed a variational reasoning method to estimate explicit user preferences during the dialogue.
@@
Privacy protection has received more and more attention in dialogue systems [papernot2016semi, henderson2018ethical]. For many task-oriented dialogue systems, it is necessary to notice that we are using the same dialogue assistant. Recent studies on membership inference attacks have confirmed that privacy information in training data for sequence-to-sequence generative models and pre-trained language models can be attacked [hisamoto2020membership, liu2021encodermi]. As we discussed in Section 7.2, most existing e-commerce dialogue system models are designed based on sequence-to-sequence generative neural networks and pre-trained language models. By learning through interactions and communications, a dialogue assistant can inadvertently and implicitly store some sensitive information. Hence consumers’ privacy information may get obtained by attackers through membership inference attacks. To address this problem, developing privacy-aware dialogue systems will attract increasing attention in the future.
@@
The evaluation of e-commerce dialogue systems is a crucial part of the development process. Recent studies on evaluating dialogue systems are either through offline evaluation or human evaluation [Lamel2000TheLA, Jurccek2011RealUE]. Offline evaluation is often limited to single-turn assessments, while human evaluation is intrusive, time-intensive, and does not scale [Deriu2020SurveyOE]. User simulators have been applied to exhaustively enumerate user goals to generate human-like conversations for simulated evaluation [Zhang2020RecentAA]. However, user simulators as evaluation methods for e-commerce dialogue systems are still under-explored. Employing existing simulators suffers from limited realism and evaluation capability, whereas there is a lack of automatic methods to assess user simulators’ realism and evaluation capabilities [Balog2021Sim4IRTS]. Moreover, evaluation metrics of e-commerce aspects are still unexploited in existing e-commerce dialogue agents. Evaluation metrics in e-commerce search and recommendation scenarios can provide more insights for future work.
@@
We briefly summarize the main topics presented in this survey in Section 8.1. In Section 8.2, we describe our outlook on future developments.

### v3

@@ 7.2.3 Knowledge-grounded dialogue systems
Although answering inquiries is essential for dialogue systems, especially for task-oriented dialogue systems, it is still far behind a natural knowledge-grounded dialogue system, which should be able to understand the facts involved in the current dialogue session (so-called fact matching) and diffuse them to other similar entities for knowledge-based dialogues (i.e., entity diffusion):
@@
    Fact matching: In dialogue systems, matching utterances to exact facts is much harder than answering explicit factoid inquiries. Though some utterances, whose subjects and relations can be easily recognized, are fact-related inquiries, the subjects and relations are often elusive, leading to challenges when matching exact facts. Table 7.1 shows an example, with items 1 and 2 talking about the film “Titanic.” Unlike item 1, which is a typical question-answering conversation, item 2 is a knowledge-related chit-chat without any explicit relation. It is difficult to define the exact fact match for item 2.
@@
    Entity diffusion: Conversations usually drift from one entity to another. In Table 7.1, the utterances in items 3 and 4 are about the entity “Titanic.” However, the entities in the responses are other similar films. Current knowledge triplets rarely capture such entity diffusion relations. The response in item 3 shows that the two entities, “Titanic” and “Waterloo Bridge,” are relevant through “love stories.” Item 4 suggests another similar shipwreck film “Titanic.”
@@ 4 A: Is there anything like the Titanic?
Knowledge-grounded dialogue systems address the aforementioned challenges. Work on knowledge-grounded dialogue systems can be categorized into two groups. Methods in the first group use structured knowledge (given knowledge graphs) [wu2020diverse, zhou-etal-2020-kdconv, wang2020improving, wu2020improving, xu2020knowledge, xu2020conversational, jung2020attnio, tuan2019dykgchat, wu2019proactive, moon2019opendialkg, zhou2018commonsense, liu2018knowledge]. Methods in the second group focus on using unstructured knowledge, such as document-based unstructured knowledge (given a whole document, e.g., a Wikipedia article) [chuanmeng2020refnet, ma2020survey, ma2020compare, tian-etal-2020-response, ren2019thinking, gopalakrishnan2019topical, zekangli2018incremental, qin-etal-2019-conversing, moghe2018towards, zhou2018dataset, parthasarathi2018extending] or piece-based unstructured knowledge (given some separate pieces of knowledge, e.g., Foursquare tips) [ghazvininejad2018knowledge, dinan2018wizard, meng2020dukenet, kim2020sequential, Lian2019Learning, zheng2020difference, chen2020bridging, zheng2020approximation, zhao2020knowledge, zheng2019enhancing, lin-etal-2020-generating]. There are key research directions for both groups: (i) improving knowledge selection [kim2020sequential]; (ii) improving knowledge-aware response generation [zhao2020knowledge] or response selection [young2018augmenting, zhao2019document, hua2020learning, sun2020history]; that is, given the chosen knowledge, how to better generate a response token by token or select a response from pre-defined response candidates; (iii) using multiple knowledge modalities [liu2019knowledge, zhao2020multiple]; that is, how to use structured, unstructured, and even other types of knowledge simultaneously; (iv) overcoming data scarcity [zhao2019low, li2020zero].
@@
The neural knowledge diffusion model introduces knowledge into the dialogue generation. This method can match the relevant facts for the input utterance and diffuse them to similar entities [liu2018knowledge]. Early studies on knowledge selection in knowledge-grounded dialogue systems calculate the weight of each piece of knowledge and obtain a weighted sum of their representations [ghazvininejad2018knowledge, zheng2019enhancing, lin-etal-2020-generating, zheng2020approximation]. bordes2016learning employ memory networks to address restaurant reservations, using a small number of keywords to handle entity types in a knowledge base (cuisine type, location, price range, party size, rating, phone number, and address). ghazvininejad2018knowledge adapt it to memorize relevant, grounded facts for a neural conversation model. The hierarchical variational memory network (HVMN) adds hierarchical structure and a variational memory network into a neural encoder-decoder network for non-task-oriented dialogue generation [ziwwsdm1].
@@
Several recent studies on knowledge selection focus on calculating a weight on each piece of knowledge and then directly sampling the amount of knowledge with the highest weight. Specifically, dinan2018wizard design the TMemNet model that uses context to predict a distribution over pieces of knowledge and then only sample one of them into a decoder. They also introduce a knowledge selection loss to supervise knowledge selection during training. Lian2019Learning describe PostKS, which uses a context to predict a prior distribution over pieces of knowledge. During training, the prior distribution is supervised by a posterior distribution predicted by the context and the corresponding response. Similar to PostKS, zheng2020approximation use a context and a piece of knowledge retrieved by the context to predict a distribution over fragments of knowledge, where the probability of the amount of knowledge retrieved by the corresponding response is maximized during training. The former distills a context containing multiple utterances at different turns into a vector that is used to match with the representation of a piece of knowledge to get a score, while the latter matches every utterance in a context with a piece of knowledge to get matching features that are aggregated to get a score. A piece of knowledge is chosen based on the score list for all pieces of knowledge. kim2020sequential propose a sequential knowledge transformer (SKT), which jointly uses previously selected knowledge and context to facilitate knowledge selection. chen2020bridging upgrade SKT by adding the knowledge distillation-based training strategy to improve knowledge selection. zheng2020difference detail a method that introduces the difference information between the previously selected knowledge and the current pieces of candidate knowledge to facilitate knowledge selection. meng2020dukenet design DukeNet, which regards tracking the previously selected knowledge and selecting the current knowledge as dual tasks within a dual learning paradigm [qin2020dual]. zhao2020knowledge describe a method, RLKS, where the selected knowledge is sent to a decoder to generate a response that would be compared with the ground truth response to give feedback to further supervised knowledge selection. meng2021initiative introduce a mixed-initiative knowledge selection method for knowledge-grounded conversations that explicitly distinguishes between user-initiative and system-initiative knowledge selection at each conversation turn to improve the performance of knowledge selection. sun2021conversations find that the amount of knowledge available in different languages is highly unbalanced. Hence, the authors address cross-lingual knowledge grounded conversations with a self-distillation knowledge selection and curriculum learning.
@@
More recent years have witnessed the rapid development of pre-trained language models in open-domain dialogue systems. Large pre-trained language models can store knowledge into their parameters during pre-training and can generate informative responses in conversations [Zhao2020ArePL]. Petroni2019LanguageMA have shown that pre-trained language models can serve as knowledge bases for downstream tasks (e.g., question-answering [Roberts2020HowMK]). On this basis, Zhao2020ArePL have shown that pre-trained language models can ground open-domain dialogues using their implicit knowledge. madotto2020learning embed knowledge bases into model’s parameters for end-to-end task-oriented dialogues. Roller2021RecipesFB fine-tune pre-trained language models on knowledge-grounded conversational data. Cui2021KnowledgeEF describe knowledge-enhanced fine-tuning methods to handle unseen entities. Xu2021RetrievalFreeKD propose a topic-aware adapter to adapt pre-trained language models in knowledge-grounded dialogues. Liu2022MultiStagePF introduce a multi-stage prompting approach for triggering knowledge in pre-trained language models. Wu2022LexicalKI design lexical knowledge internalization to integrate token-level knowledge into the model’s parameters. The problem of hallucination is becoming more and more challenging. sun2023contrastive optimize an implicit knowledge eliciting process, i.e., they reduce hallucination of pre-trained language models in knowledge-grounded dialogues through a contrastive learning framework.
@@ 7.2.4 Empathetic dialogue generation
Several approaches to data-driven open-domain dialogue generation generate emotional responses based on a manually specified label to control the dynamic content of the target output [zhou2018emotional, lis18, zhou2018mojitalk, huang2018automatic, wei2019emotion, colombo2019affect, shenf20].
Unlike emotional dialogue generation, the study of empathetic dialogue generation avoids an additional step of determining which emotion type to respond to explicitly [skowron2013affect]. Several studies [Rashkin18, zhong2019affect, Shin19, rashkin2019towards, santhanam19, linmsxf19, lin2020caire, zhong2020, MajumderHPLGGMP20, li2020empdg] have attempted to make dialogue models more empathetic. rashkin2019towards combine models in different ways to produce empathetic responses. linmsxf19 softly combine the possible emotional responses from several separate experts. MajumderHPLGGMP20 consider polarity-based emotion clusters and emotional mimicry. li2020empdg propose a multi-resolution adversarial framework that considers multi-granularity emotion factors and users’ feedback. li2020empathetic investigate how to use external knowledge to explicitly improve the emotional understanding and expression in the task of empathetic dialogue generation. sabour2022cem focus on two aspects in empathetic dialogue generation: affection and cognition. The authors propose a method with various commonsense reasoning to improve understanding of interlocutors’ situations and feelings.
Besides advances in empathetic dialogue models, the emergence of new emotion-labeled dialogue corpora has also contributed to this research field [li2017dailydialog, hsuckhk18, rashkin2019towards]. rashkin2019towards consider a rich and well-balanced set of emotions and release a dataset, EmpatheticDialogues, where a listener responds to a speaker in an emotional situation in an empathetic way.
@@ 7.3 Emerging directions
This section describes recent emerging directions on question-answering and dialogue systems in e-commerce. These emerging directions in QA and dialogue systems can be divided into five perspectives: safety, ethics, interpretability, privacy, and evaluation.
@@
As discussed in Section 7.1, e-commerce QA agents aim to answer questions based on large volumes of reviews. However, reviews may not answer these questions as they may not contain any relevant answers for the question, or a query may be poorly phrased and therefore require additional clarification. Moreover, untruthful comments and spam are widely observed in e-commerce reviews [carmel2018product]. mihaylova2019semeval investigate the fact-checking problem in a QA scenario with a system to classify the veracity of answers. zhang2020answerfact release a large-scale fact-checking dataset called AnswerFact for investigating the answer veracity in e-commerce QA. estes2022fact develop a high-speed fact-verification system that has a very high false statement recall and very high true statement precision to product question-answering. However, the authors still apply a rule-based method in their system, and find that pre-trained language models are unable to perform fact-checking well on structured catalog data. As limitations such as poor generalization exist in rule-based methods, how to optimize pre-trained language models in fact-checking for product question-answering still needs more attention in future research.
@@
Ethical challenges in dialogue systems are attracting significant amounts of attention in recent years. Currently, most dialogue systems are developed from scratch with large corpora or fine-tuned through pre-trained language models. Large-scale datasets collected from the open internet have been applied during model training. However, offensive and malevolent content can be observed in the data [si2022so]. To avoid being unintentionally offensive or harming the user, studies have been performed to detect toxic speech around, e.g., religion, race, and violence [tripathi2019detecting, dinan2020queens, zhang2021human, kann2022open, si2022so]. zhang2021human propose a human-machine collaborative evaluation framework for reliable toxic speech detection in dialogue systems. si2022so study toxic speech in open-domain dialogue systems to reveal that specific kinds of “non-toxic” queries are able to trigger an open-domain conversational assistant to output toxic responses. However, how to respond when these malevolence topics are being identified is still an open question [kann2022open]. As more and more e-commerce dialogue systems have also been trained based on pre-trained language models, ethical challenges will need to be tackled in future research.
@@
Poor explainability is a challenging problem for most e-commerce QA approaches. Most e-commerce QA approaches apply end-to-end semantic matching methodologies, which tend to be black-box and directly output a matching score for each question answer pair. zhao2019riker address the explainable QA problem through a hybrid retrieval-based framework. The authors employ a bidirectional recurrent neural network in the internal word representation stage and apply a keyword-aware retrieval method during the second stage. In contrast, the tf-idf ranking function naturally exhibits much better interpretability owing to its transparency and intuitiveness. Conversational recommendation is another typical application of e-commerce dialogue systems [sun2016conversational, mangili2020bayesian]. Most approaches neglect explainability when learning recommendation actions. However, chen2020towards propose an incremental multi-task learning framework using user feedback for the task of explainable conversational recommendation. By considering user preferences as latent variables in a variational Bayesian manner, ren2022variational employ a method to estimate explicit user preferences during the dialogue.
@@
Privacy protection has received more and more attention in dialogue systems [papernot2016semi, henderson2018ethical]. For many task-oriented dialogue systems, it is necessary to notice that we are using the same dialogue assistant. Recent studies on membership inference attacks have confirmed that privacy information in training data for sequence-to-sequence generative models and pre-trained language models can be attacked [hisamoto2020membership, liu2021encodermi]. As we discussed in Section 7.2, most e-commerce dialogue system models are designed based on sequence-to-sequence generative neural networks and pre-trained language models. By learning through interactions and communications, a dialogue assistant can inadvertently and implicitly store sensitive information. Hence, consumers’ privacy information may get obtained by attackers through membership inference attacks. To address this problem, developing privacy-aware dialogue systems is likely to attract increased attention in the future.
@@
The evaluation of e-commerce dialogue systems is a crucial part of the development process. Recent studies on evaluating dialogue systems are either through offline evaluation or human evaluation [Lamel2000TheLA, Jurccek2011RealUE]. Offline evaluation is often limited to single-turn assessments, while human evaluation is intrusive, time-intensive, and does not scale [Deriu2020SurveyOE]. User simulators have been applied to exhaustively enumerate user goals to generate human-like conversations for simulated evaluation [Zhang2020RecentAA]. However, user simulators as evaluation methods for e-commerce dialogue systems are still under-explored. Today’s simulators suffer from limited realism and evaluation capabilities [Balog2021Sim4IRTS]. Moreover, evaluation metrics that specifically target e-commerce aspects are still underexplored and underexploited in today’s e-commerce dialogue agents. Dedicated evaluation metrics in e-commerce search and recommendation scenarios will likely help to advance progress.
@@
We summarize the main topics presented in this survey in Section 8.1. In Section 8.2, we describe our outlook on future developments.